# ReconX: Reconstruct Any Scene from Sparse Views with Video Diffusion Model

### v2

@@ Abstract
Advancements in 3D scene reconstruction have transformed 2D images from the real world into 3D models, producing realistic 3D results from hundreds of input photos. Despite great success in dense-view reconstruction scenarios, rendering a detailed scene from sparse views is still an ill-posed optimization problem, often resulting in artifacts and distortions in unseen areas. In this paper, we propose ReconX, a novel 3D scene reconstruction paradigm that reframes the ambiguous reconstruction problem as a temporal generation task. The key insight is to unleash the strong generative prior of large pre-trained video diffusion models for sparse-view reconstruction. Nevertheless, it is challenging to preserve 3D view consistency when directly generating video frames from pre-trained models. To address this issue, given limited input views, the proposed ReconX first constructs a global point cloud and encodes it into a contextual space as the 3D structure condition. Guided by the condition, the video diffusion model then synthesizes video frames that are detail-preserved and exhibit a high degree of 3D consistency, ensuring the coherence of the scene from various perspectives. Finally, we recover the 3D scene from the generated video through a confidence-aware 3D Gaussian Splatting optimization scheme. Extensive experiments on various real-world datasets show the superiority of ReconX over state-of-the-art methods in terms of quality and generalizability. Project Page: https://liuff19.github.io/ReconX.
1 Introduction
@@
With the rapid development of photogrammetry techniques such as NeRF (Mildenhall et al., 2020) and 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023), 3D reconstruction has become a popular research topic in recent years, finding various applications from virtual reality (Dalal et al., 2024) to autonomous navigation (Adamkiewicz et al., 2022) and beyond (Martin-Brualla et al., 2021b; Liu et al., 2024a; Wu et al., 2024a; Charatan et al., 2024). However, sparse-view reconstruction is an ill-posed problem (Gao et al., 2024; Yu et al., 2021) since it involves recovering a complex 3D structure from limited viewpoint information (i.e., even as few as two images) that may correspond to multiple solutions. This uncertain process requires additional assumptions and constraints to yield a viable solution.
@@
Recently, powered by the efficient and expressive 3DGS (Kerbl et al., 2023) with fast rendering speed and high quality, several feed-forward Gaussian Splatting methods (Charatan et al., 2024; Szymanowicz et al., 2024b; Chen et al., 2024a) have been proposed to explore 3D scene reconstruction from sparse view images. Although they can achieve promising interpolation results by learning scene-prior knowledge from feature extraction modules (e.g., epipolar transformer (Charatan et al., 2024)), insufficient captures of the scene still lead to an ill-posed optimization problem (Wu et al., 2024b). As a result, they often suffer from severe artifact and implausible imagery issues when rendering the 3D scene from novel viewpoints, especially in unseen areas.
To address the limitations, we propose ReconX, a novel 3D scene reconstruction paradigm that reformulates the inherently ambiguous reconstruction problem as a generation problem. Our key insight is to unleash the strong generative prior of pre-trained large video diffusion models (Blattmann et al., 2023a; b; Xing et al., 2023) to create more observations for the downstream reconstruction task. Despite the capability to synthesize video clips featuring plausible 3D structures (Gao et al., 2024), recovering a high-quality 3D scene from current video diffusion models is still challenging, due to the poor 3D view consistency across generated 2D frames. Grounded by theoretical analysis, we explore the potential of incorporating 3D structure condition into the video generative process, which bridges the gap between the under-determined 3D creation problem and the fully-observed 3D reconstruction setting. Specifically, given sparse images, we first build a global point cloud through a pose-free stereo reconstruction method. Then we encode it into a rich context representation space as the 3D condition in cross-attention layers, which guides the video diffusion model to synthesize detail-preserved frames with 3D consistent novel observations of the scene. Finally, we reconstruct the 3D scene from the generated video through Gaussian Splatting with a 3D confidence-aware and robust scene optimization scheme, which further deblurs the uncertainty in video frames effectively. Extensive experiments verify the efficacy of our framework and show that ReconX outperforms existing methods for high quality and generalizability, revealing the great potential to craft intricate 3D worlds from video diffusion models. The overview and examples of reconstructions are shown in Figure 1.
@@ 2 Related Work
Sparse-view reconstruction. NeRF and 3DGS typically demand hundreds of input images and rely on the multi-view stereo reconstruction (MVS) approach (e.g., COLMAP (Schönberger & Frahm, 2016)) to estimate the camera parameters. To address the issue of low-quality 3D reconstruction caused by sparse views, PixelNeRF (Yu et al., 2021) proposes using convolutional neural networks to extract features from the input context. Moreover, FreeNeRF (Yang et al., 2023) adopts the frequency and density regularized strategies to alleviate the artifacts caused by insufficient inputs without any additional cost. To mitigate the overfitting to input sparse views in 3DGS, FSGS (Zhu et al., 2023) and SparseGS (Xiong et al., 2023) employ a depth estimator to regularize the optimization process. However, these methods all require known camera intrinsics and extrinsics, which is not practical in real-world scenario. Benefiting from the existing powerful 3D reconstruction model (i.e., DUSt3R (Wang et al., 2024a)), InstantSplat (Fan et al., 2024) is able to acquire accurate camera parameters and initial 3D representations from unposed sparse-view inputs, leading to the efficient and high-quality 3D reconstruction.
@@
Regression model for generalizable view synthesis. While NeRF and 3DGS are optimized per-scene, a line of research aims to train feed-forward models that output a 3D representation directly from a few input images, bypassing the need for time-consuming optimization. Splatter image (Szymanowicz et al., 2024b) performs an efficient feed-forward manner for monocular 3D object reconstruction by predicting a 3D Gaussian for each image pixel. Meanwhile, pixelSplat (Charatan et al., 2024) proposes predicting the scene-level 3DGS from the image pairs, using the epipolar transformer to better extract scene features. Following that, MVsplat (Chen et al., 2024a) introduces the cost volume and depth refinements to produce a clean and high-quality 3D Gaussians in a faster way. LatentSplat (Wewer et al., 2024) encodes the variational 3D Gaussians and utilizes a discriminator to synthesize more realistic images. To reconstruct a complete scene from a single image, Flash3D (Szymanowicz et al., 2024a) adopts a hierarchical 3DGS learning policy and depth constraint to achieve high-quality interpolation and extrapolation view synthesis. Although these methods leverage the 3D data priors, they are limited by the scarcity and diversity of 3D data. Consequently, these methods struggle to achieve high-quality renderings in unseen areas, especially when out-of-distribution (OOD) data is used as input.
Generative models for 3D reconstruction. Constructing comprehensive 3D scenes from limited observations demands generating 3D content, particularly for unseen areas. Earlier studies distill the knowledge in the pre-trained text-to-image diffusion models (Rombach et al., 2022; Saharia et al., 2022; Ramesh et al., 2022) into a coherent 3D model. Specifically, the Score Distillation Sampling (SDS) technique (Wu et al., 2024b; Lin et al., 2023; Liu et al., 2024c; Wang et al., 2024b) is adopted to synthesize a 3D object from the text prompt. To enhance the 3D consistency, several approaches (Wu et al., 2024a; Shi et al., 2023; Liu et al., 2023) inject the camera information into diffusion models, providing strong multi-view priors. Furthermore, ZeroNVS (Sargent et al., 2023) and CAT3D (Gao et al., 2024) extend the multi-view diffusion to the scene level generation. GeNVS (Chan et al., 2023) embeds a 3D feature field into the diffusion model to enhance the novel view synthesis ability. More recently, video diffusion models (Blattmann et al., 2023a; Xing et al., 2023) have shown an impressive ability to produce realistic videos and are believed to implicitly understand 3D structures (Liu et al., 2024b). SV3D (Voleti et al., 2024) and V3D (Chen et al., 2024b) explore fine-tuning the pre-trained video diffusion model for 3D object generation. Meanwhile, MotionCtrl (Wang et al., 2024c) and CameraCtrl (He et al., 2024) achieve scene-level controllable video generation from a single image by explicitly injecting the camera pose into video diffusion models. However, they cannot work for the unconstrained sparse-view 3D scene reconstruction, which requires strong 3D consistency.
3 Motivation for ReconX
In this paper, we focus on the fundamental problem of 3D scene reconstruction and novel view synthesis (NVS) from very sparse view (e.g., as few as two) images. Most existing works (Chen et al., 2024a; Yu et al., 2021; Charatan et al., 2024; Szymanowicz et al., 2024a) utilize 3D prior and geometric constraints (e.g., depth, normal, cost volume) to fill the gap between observed and novel regions in sparse-view 3D reconstruction. Although capable of producing highly realistic images from the given viewpoints, these methods often struggle to generate high-quality images in areas not visible from the input perspectives due to the inherent problem of insufficient viewpoints and the resulting instability in the reconstruction process. To address this issue, a natural idea is to create more observations to convert the under-determined 3D creation problem into a fully constrained 3D reconstruction setting. Recently, video generative models have shown promise for synthesizing video clips featuring 3D structures (Voleti et al., 2024; Blattmann et al., 2023a; Xing et al., 2023). This inspires us to unleash the strong generative prior of large pre-trained video diffusion models to create temporal consistent video frames for sparse-view reconstruction. Nevertheless, it is non-trivial as the main challenge lies in poor 3D view consistency among video frames, which significantly limits the downstream 3DGS training process. To achieve 3D consistency within video generation, we first analyze the video diffusion modeling from a 3D distributional view. Let $\bm{x}$ be the set of rendering 2D images from any 3D scene in the world, $q(\bm{x})$ be the distribution of the rendering data $\bm{x}$, and our goal is to minimize the divergence $\mathcal{D}$:
$\min_{\bm{\theta}\in\Theta,\psi\in\Psi}\mathcal{D}\left(q(\bm{x})\|p_{\bm{\theta},\psi}(\bm{x})\right),$ (1)
where $p_{\bm{\theta},\psi}$ is a diffusion model parameterized by $\bm{\theta}\in\Theta$ (the parameters in the backbone) and $\psi\in\Psi$ (any embedding function shared by all data). The vanilla video diffusion model (Xing et al., 2023) chooses a CLIP (Radford et al., 2021) model $g$ to add an image-based condition (i.e., $\psi=g$). However, in sparse-view 3D reconstruction, only conditioning on 2D images cannot provide sufficient condition for approximating $q(\bm{x})$ (Charatan et al., 2024; Chen et al., 2024a; Wu et al., 2024b). Motivated by this, we explore the potential of incorporating the native 3D prior (denoted by $\mathcal{F}$) to find an optimal solution in Equation 1 and derive a theoretical formulation for our analysis in Proposition 1.
Proposition 1.
Let $\bm{\theta}^{*},\psi^{*}=g^{*}$ be the optimal solution of the solely image-based conditional diffusion scheme and $\bm{\tilde{\theta}}^{*},\tilde{\psi}^{*}=\{{g}^{*},\mathcal{F}^{*}\}$ be the optimal solution of the diffusion scheme with a native 3D prior. Suppose the divergence $\mathcal{D}$ is convex and the embedding function space $\Psi$ includes all measurable functions, then we have $\mathcal{D}(q\left(\bm{x}\right)\|p_{\bm{\tilde{\theta}}^{*},\tilde{\psi}^{*}}(\bm{x}))<\mathcal{D}\left(q\left(\bm{x}\right)\|p_{\bm{\theta}^{*},\psi^{*}}(\bm{x})\right)$. (Proof in our Appendix)
Towards this end, we reformulate the inherently ambiguous reconstruction problem as a generation problem by incorporating a 3D native structure condition into the diffusion process. More preliminaries can be found in our Appendix.
4 Method
@@
4.1 Overview of ReconX

### v3

@@ Abstract
Advancements in 3D scene reconstruction have transformed 2D images from the real world into 3D models, producing realistic 3D results from hundreds of input photos. Despite great success in dense-view reconstruction scenarios, rendering a detailed scene from sparse views is still an ill-posed optimization problem, often resulting in artifacts and distortions in unseen areas. In this paper, we propose ReconX, a novel 3D scene reconstruction paradigm that reframes the ambiguous reconstruction problem as a temporal generation task. The key insight is to unleash the strong generative prior of large pre-trained video diffusion models for sparse-view reconstruction. Nevertheless, it is challenging to preserve 3D view consistency when directly generating video frames from pre-trained models. To address this issue, given limited input views, the proposed ReconX first constructs a global point cloud and encodes it into a contextual space as the 3D structure condition. Guided by the condition, the video diffusion model then synthesizes video frames that are detail-preserved and exhibit a high degree of 3D consistency, ensuring the coherence of the scene from various perspectives. Finally, we recover the 3D scene from the generated video through a confidence-aware 3D Gaussian Splatting optimization scheme. Extensive experiments on various real-world datasets show the superiority of ReconX over state-of-the-art methods in terms of quality and generalizability.
Index Terms:
Sparse-view Reconstruction, Video Diffusion, Gaussian Splatting
I Introduction
@@
With the rapid development of photogrammetry techniques such as NeRF [1] and 3D Gaussian Splatting (3DGS) [2], 3D reconstruction has become a popular research topic in recent years, finding various applications from virtual reality [3] to autonomous navigation [4] and beyond [5, 6, 7, 8, 9]. However, sparse-view reconstruction is an ill-posed problem [10, 11] since it involves recovering a complex 3D structure from limited viewpoint information (i.e., even as few as two images) that may correspond to multiple solutions. This uncertain process requires additional assumptions and constraints to yield a viable solution.
@@
Recently, powered by the efficient and expressive 3DGS [2] with fast rendering speed and high quality, several feed-forward Gaussian Splatting methods [12, 13, 14] have been proposed to explore 3D scene reconstruction from sparse view images. Although they can achieve promising interpolation results by learning scene-prior knowledge from feature extraction modules (e.g., epipolar transformer [12]), insufficient captures of the scene still lead to an ill-posed optimization problem [15]. As a result, they often suffer from severe artifact and implausible imagery issues when rendering the 3D scene from novel viewpoints, especially in unseen areas.
@@
To address the limitations, we propose ReconX, a novel 3D scene reconstruction paradigm that reformulates the inherently ambiguous reconstruction problem as a generation problem. Our key insight is to unleash the strong generative prior of pre-trained large video diffusion models [16, 17, 18] to create more observations for the downstream reconstruction task. Despite the capability to synthesize video clips featuring plausible 3D structures [10], recovering a high-quality 3D scene from current video diffusion models is still challenging, due to the poor 3D view consistency across generated 2D frames. Grounded by theoretical analysis, we explore the potential of incorporating 3D structure condition into the video generative process, which bridges the gap between the under-determined 3D creation problem and the fully-observed 3D reconstruction setting. Specifically, given sparse images, we first build a global point cloud through a pose-free stereo reconstruction method. Then we encode it into a rich context representation space as the 3D condition in cross-attention layers, which guides the video diffusion model to synthesize detail-preserved frames with 3D consistent novel observations of the scene. Finally, we reconstruct the 3D scene from the generated video through Gaussian Splatting with a 3D confidence-aware and robust scene optimization scheme, which further deblurs the uncertainty in video frames effectively. Extensive experiments verify the efficacy of our framework and show that ReconX outperforms existing methods for high quality and generalizability, revealing the great potential to craft intricate 3D worlds from video diffusion models. The overview and examples of reconstructions are shown in Fig. 1.
@@ II Related Work
Sparse-view reconstruction. NeRF [1] and 3DGS [2] typically demand hundreds of input images and rely on the multi-view stereo reconstruction (MVS) approach (e.g., COLMAP [19]) to estimate the camera parameters. To address the issue of low-quality 3D reconstruction caused by sparse views, PixelNeRF [11] proposes using convolutional neural networks to extract features from the input context. Moreover, FreeNeRF [20] adopts the frequency and density regularized strategies to alleviate the artifacts caused by insufficient inputs without any additional cost. To mitigate the overfitting to input sparse views in 3DGS, FSGS [21] and SparseGS [22] employ a depth estimator to regularize the optimization process. However, these methods all require known camera intrinsics and extrinsics, which is not practical in real-world scenario. Benefiting from the existing powerful 3D reconstruction model (i.e., DUSt3R [23]), InstantSplat [24] is able to acquire accurate camera parameters and initial 3D representations from unposed sparse-view inputs, leading to efficient and high-quality 3D scene reconstruction.
@@
Regression model for generalizable view synthesis. While 3D reconstruction methods like NeRF and 3DGS are optimized per-scene [25, 26, 27, 28, 29, 30], a line of research aims to train feed-forward models that output a 3D representation directly from a few input images, bypassing the need for time-consuming optimization. Splatter image [13] performs an efficient feed-forward manner for monocular 3D object reconstruction by predicting a 3D Gaussian for each image pixel. Meanwhile, pixelSplat [12] proposes predicting the scene-level 3DGS from the image pairs, using the epipolar transformer to better extract scene features. Following that, MVsplat [14] introduces the cost volume and depth refinements to produce a clean and high-quality 3D Gaussians in a faster way. LatentSplat [31] encodes the variational 3D Gaussians and utilizes a discriminator to synthesize more realistic images. To reconstruct a complete scene from a single image, Flash3D [32] adopts a hierarchical 3DGS learning policy and depth constraint to achieve high-quality interpolation and extrapolation view synthesis. Although these methods leverage the 3D data priors, they are limited by the scarcity and diversity of 3D data. Consequently, these methods struggle to achieve high-quality renderings in unseen areas, especially when out-of-distribution (OOD) data is used as input.
@@
Generative models for 3D reconstruction. Constructing comprehensive 3D scenes from limited observations demands generating 3D content, particularly for unseen areas. Earlier studies distill the knowledge in the pre-trained text-to-image diffusion models [33, 34, 35, 36] into a coherent 3D model. Specifically, the Score Distillation Sampling (SDS) technique [15, 37, 38, 39] is adopted to synthesize a 3D object from the text prompt. To enhance the 3D consistency, several approaches [8, 40, 41] inject the camera information into diffusion models, providing strong multi-view priors. Furthermore, ZeroNVS [42] and CAT3D [10] extend the multi-view diffusion to the scene level generation. GeNVS [43] embeds a 3D feature field into the diffusion model to enhance the novel view synthesis ability. More recently, video diffusion models [16, 18] have shown an impressive ability to produce realistic videos and are believed to implicitly understand 3D structures [44]. SV3D [45] and V3D [46] explore fine-tuning the pre-trained video diffusion model for 3D object generation. Meanwhile, MotionCtrl [47] and CameraCtrl [48] achieve scene-level controllable video generation from a single image by explicitly injecting the camera pose into video diffusion models. However, they suffer from performance degradation in the unconstrained sparse-view 3D scene reconstruction, which requires strong 3D consistency.
III Preliminaries
Video Diffusion Models. Diffusion models [49, 50] have emerged as the cutting-edge paradigm to generate high-quality videos. These models learn the underlying data distribution by adding and removing noise on the clean data. The forward process aims to transform a clean data sample $\boldsymbol{x}_{0}\sim p(\boldsymbol{x})$ to a pure Gaussian noise $\boldsymbol{x}_{T}\sim\mathcal{N}(0,I)$, following the process:
$\boldsymbol{x}_{t}=\sqrt{\bar{\alpha}_{t}}\boldsymbol{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,\quad\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{1}),$ (1)
where $\boldsymbol{x}_{t}$ and $\bar{\alpha}_{t}$ denotes the noisy data and noise strength at the timestep $t$. The denoising neural network $\epsilon_{\theta}$ is trained to predict the noises added in the forward process, which is achieved by the MSE loss:
$\mathcal{L}=\mathbb{E}_{\boldsymbol{x}\sim p,\epsilon\sim\mathcal{N}(0,I),c,t}\left[\left\|\epsilon-\epsilon_{\theta}\left(\boldsymbol{x}_{t},t,c\right)\right\|_{2}^{2}\right],$ (2)
where $c$ represents the embeddings of conditions like text or image prompt. For the video diffusion models, Latent Diffusion Models (LDMs) [33], which compress images into the latent space, are commonly employed to mitigate the computation complexity while maintaining competitive performance.
3D Gaussian Splatting. 3DGS [2] represents a scene explicitly by utilizing a set of 3D Gaussian spheres, achieving a fast and high-quality rendering. A 3D Gaussian is modeled by a position vector $\boldsymbol{\mu}\in\mathbb{R}^{3}$, a covariance matrix $\boldsymbol{\Sigma}\in\mathbb{R}^{3\times 3}$, an opacity $\alpha\in\mathbb{R}$, and spherical harmonics (SH) coefficient $\boldsymbol{c}\in\mathbb{R}^{k}$ [51]. Moreover, the Gaussian distribution is formulated as the following:
$G(x)=e^{-\frac{1}{2}(x-\boldsymbol{\mu})^{T}\boldsymbol{\Sigma}^{-1}(x-\boldsymbol{\mu)}},$ (3)
where $\boldsymbol{\Sigma}=\boldsymbol{R}\boldsymbol{S}\boldsymbol{S}^{T}\boldsymbol{R}^{T}$, $\boldsymbol{S}$ denotes the scaling matrix and $\boldsymbol{R}$ is the rotation matrix.
In the rendering stage, the 3D Gaussian spheres are transformed into 2D camera planes through rasterization [52]. Specifically, given the perspective transformation matrix $\boldsymbol{W}$ and Jacobin of the projection matrix $\boldsymbol{J}$, the 2D covariance matrix in the camera space is computed as
$\boldsymbol{\Sigma}^{{}^{\prime}}=\boldsymbol{J}\boldsymbol{W}\boldsymbol{\Sigma}\boldsymbol{W}^{T}\boldsymbol{J}^{T}.$ (4)
End-to-end Dense Unconstrained Stereo. DUSt3R [23] is a new model to predict a dense and accurate 3D scene representation solely from image pairs without any prior information about the scene. Given two unposed images $\{\boldsymbol{I}_{1},\boldsymbol{I}_{2}\}$, this end-to-end model is trained to estimate the point maps $\{P_{1,1},P_{2,1}\}$ and confidence maps $\{\mathcal{C}_{1,1},\mathcal{C}_{2,1}\}$, which can be utilized to recover the camera parameters and dense point cloud. The training procedure for view $v\in\{1,2\}$ is formulated as a regression loss:
IV Method
IV-A Motivation for ReconX
In this paper, we focus on the fundamental problem of 3D scene reconstruction and novel view synthesis (NVS) from very sparse view (e.g., as few as two) images. Most existing works [14, 11, 12, 32] utilize 3D prior and geometric constraints (e.g., depth, normal, cost volume) to fill the gap between observed and novel regions in sparse-view 3D reconstruction. Although capable of producing highly realistic images from the given viewpoints, these methods often struggle to generate high-quality images in areas not visible from the input perspectives due to the inherent problem of insufficient viewpoints and the resulting instability in the reconstruction process. To address this issue, a natural idea is to create more observations to convert the under-determined 3D creation problem into a fully constrained 3D reconstruction setting. Recently, video generative models have shown promise for synthesizing video clips featuring 3D structures [45, 16, 18]. This inspires us to unleash the strong generative prior of large pre-trained video diffusion models to create temporal consistent video frames for sparse-view reconstruction. Nevertheless, it is non-trivial as the main challenge lies in poor 3D view consistency among video frames, which significantly limits the downstream 3DGS training process. To achieve 3D consistency within video generation, we first analyze the video diffusion modeling from a 3D distributional view. Let $\boldsymbol{x}$ be the set of rendering 2D images from any 3D scene in the world, $q(\boldsymbol{x})$ be the distribution of the rendering data $\boldsymbol{x}$, and our goal is to minimize the divergence $\mathcal{D}$:
$\min_{\boldsymbol{\theta}\in\Theta,\psi\in\Psi}\mathcal{D}\left(q(\boldsymbol{x})\|p_{\boldsymbol{\theta},\psi}(\boldsymbol{x})\right),$ (8)
where $p_{\boldsymbol{\theta},\psi}$ is a diffusion model parameterized by $\boldsymbol{\theta}\in\Theta$ (the parameters in the backbone) and $\psi\in\Psi$ (any embedding function shared by all data). The vanilla video diffusion model [18] chooses a CLIP [53] model $g$ to add an image-based condition (i.e., $\psi=g$). However, in sparse-view 3D reconstruction, only conditioning on 2D images cannot provide sufficient condition for approximating $q(\boldsymbol{x})$ [12, 14, 15]. Motivated by this, we explore the potential of incorporating the native 3D prior (denoted by $\mathcal{F}$) to find an optimal solution in Equation 8 and derive a theoretical formulation for our analysis in Proposition 1.
Proposition 1
Let $\boldsymbol{\theta}^{*},\psi^{*}=g^{*}$ be the optimal solution of the solely image-based conditional diffusion scheme and $\boldsymbol{\tilde{\theta}}^{*},\tilde{\psi}^{*}=\{{g}^{*},\mathcal{F}^{*}\}$ be the optimal solution of the diffusion scheme with a native 3D prior. Suppose the divergence $\mathcal{D}$ is convex and the embedding function space $\Psi$ includes all measurable functions, then we have $\mathcal{D}(q\left(\boldsymbol{x}\right)\|p_{\boldsymbol{\tilde{\theta}}^{*},\tilde{\psi}^{*}}(\boldsymbol{x}))<\mathcal{D}\left(q\left(\boldsymbol{x}\right)\|p_{\boldsymbol{\theta}^{*},\psi^{*}}(\boldsymbol{x})\right)$.
Towards this end, we reformulate the inherently ambiguous reconstruction problem as a generation problem by incorporating a 3D native structure condition into the diffusion process. Detailed proof can be found in the appendix.
IV-B Overview of ReconX